OpenStudy (anonymous):

Perform the requested operation or operations. (5 points) f(x) =(x-3)/8 ; g(x) = 8x + 3, find g(f(x)).

OpenStudy (anonymous):

Does g(f(x))=x?

OpenStudy (anonymous):

probably one says multiply by 8, then add 3 the other says subtract 3, then divide by 8 so one undoes the other

jimthompson5910 (jim_thompson5910):

\[\Large \color{red}{f(x) = \frac{x-3}{8}}\] \[\Large g(x) = 8x+3\] \[\Large g(x) = 8(x)+3\] \[\Large g( \color{red}{f(x)}) = 8( \color{red}{f(x)})+3\] \[\Large g( \color{red}{f(x)}) = 8\left( \color{red}{\frac{x-3}{8}}\right)+3\] I'll let you simplify.

OpenStudy (anonymous):

So that is a yes?

jimthompson5910 (jim_thompson5910):

you tell me

OpenStudy (anonymous):

8s cancel each other out, and the the 3s do as well

jimthompson5910 (jim_thompson5910):

yes, \(\LARGE 8\left(\frac{x-3}{8}\right) + 3\) simplifies to x

jimthompson5910 (jim_thompson5910):

You'll also find that f(g(x)) = x as well. That means f(x) and g(x) are inverses of one another.

OpenStudy (anonymous):

Thanks.